Ricardo is 4 years less than twice as old as his sister. The sum of their ages is 20. How old is Ricardo?
Answer by ptaylor(2198)

Ricardo is 4 years less than twice as old as his sister. The sum of their ages is 20. How old is Ricardo?

Let x=Ricardo's sisters's age
Then 2x-4=Ricardo's age
Now we are told that the sum of their ages is 20 so our equation to solve is:
x+2x-4=20 add 4 to both sides
3x=24
x=8-----------------------------------sister's age
2x-4=2(8)-4=16-4=12-------------------------Ricardo's age
ck
Twice his sister's age is 16---4 years less that twice her age is 12--- CK
also
8+12=20
20=20



Hope this helps-----ptaylor